# Dark Energy Survey Year 6 Results: Cell-based Coadds and Metadetection Weak Lensing Shape Catalogue

### v1

@@ 2.3 Catalogue Blinding
In order to prevent accidental unblinding of cosmological information during testing, we applied a random scaling factor to the catalogue. To do this, we followed Zuntz, Sheldon et al. 2018; Gatti, Sheldon et al. 2021 to first transform ellipticity $\mathbf{e}$ into |$\mathbf{\eta}$| $\equiv$ 2arctanh(|$\mathbf{e}$|) and multiplying |$\mathbf{\eta}$| by a blinding factor $f$ (0.9<$f$<1.1). This random factor was removed once we completed the initial validation of our catalogue using the null tests detailed below.
@@ 3.1 Coadd Image Geometry
Figure 1: The cell-based coadd and DES coadd tile geometry for a corner of the DES coadd tile. The black dashed line shows the outer boundary of the 10k×$\times$10k DES coadd tile. The inner black solid line shows the unique region of the coadd tile, which is defined on the sphere in right ascension and declination. Adjacent coadd tiles overlap the same area, but share adjacent unique region boundaries. The red solid lines show the unique regions for the cell-based coadds. In the interior of the coadd tile, these regions are 100$\times$100 pixels. The red dashed line (shown only for two of the cell-based coadds along the diagonal) shows the outer boundary of each cell-based coadd. Adjacent cell-based coadds with their outer boundaries overlap by 50 pixels. At the edge of the coadd tile, the unique region for the cell-based coadd is extended all the way to the outer boundary of the coadd tile.
@@ 3.2 Cell-based Coadd Construction and Data Products
Figure 2 shows example color images of nine adjacent cells from the DES Y6 analysis on the left, the mask in the center, and the masked fraction images on the right. We applied the cell-based coadding algorithm to the $griz$ bands and have produced a $gri$ false-color image for this figure. In this figure, we have cut the 50-pixel edge buffers and stitched the inner cell images together. We can see that there are subtle shifts in the background noise as the input SE images change from cell to cell. We also see some residual striping from the interpolation applied to the input SE images. This effect is accounted for using a cut in the masked fraction around each detection when running Metadetection, as described below. The masked fraction image shows that in regions where the interpolant is applied to nearly every input SE image, and thus the output cell-based coadd image is not usable, the masked fraction approaches unity. Any source detected in or near these highly masked regions will have a large masked fraction and is excluded from use in the catalogue. The bottom left cell in Figure 2 shows the effects of the 90-degree rotation applied to the pixel masks. This rotation has not been applied to all cells in the bottom row of the images because the cells on the right are marked as being near a bright star, and we do not rotate pixel masks near bright stars since that data will be excluded anyway in our final survey mask as discussed below.
Figure 2: Left: Example false-color image of nine coadd cells (with borders indicated as white dotted lines) made from $gri$ bands. The inner 100 $\times$ 100 pixels are cut out from the whole coadd cell of 200 $\times$ 200 pixels by removing the 50 pixels buffer region around the cell. Middle: The mask image over the same region of the sky, where the grey region is excluded from the detection and further analysis. Right: The mfrac image over the same region of the sky, averaged over $riz$-bands. For each detection object, the average masked fraction is computed from the mfrac coadd image, and any objects with mfrac > 0.1 are excluded from our analysis. The interpolation and masked fraction differs for the bottom left cell relative to the two other cells on the bottom row due to the fact that 90-degree rotations are not applied to the input pixel masks near bright stars. See Sec. 3 for more details.
@@ 4.1 Shear Calibration Formalism
Metacalibration (Huff & Mandelbaum 2017; Sheldon & Huff 2017) was the first algorithm to implement this concept into practice to self-calibrate “raw” galaxy shapes. Metacalibration operates on existing object detections and includes corrections for non-detection dependent selection effects implicitly in the ensemble average in Eqn. 3. This technique deconvolves the PSF from the image, applies a small artificial shear, and then reconvolves the image with a slightly larger PSF. To this artificially sheared image, we add a noise image put through the same process but with an opposite shear applied. [^11] In detail, we rotate the noise image by 90 degrees, run an identical shearing process as the original image, and then rotate the noise image back by 90 degrees.. This correction accounts for the effect of sheared and correlated background noise on the shape measurement (Sheldon & Huff, 2017). Without it, the technique will produce catastrophic biases at the order of $\sim 10\%$. Finally, from the measured shapes on artificially sheared images, we can use a two-sided finite difference formula to compute the shear response
@@
where the subscript $i$ denotes one of the two shear components and the superscript +/- denotes the sign of the applied shear. In total, we create five images (unsheared, sheared in $+\gamma_{1},-\gamma_{1}$, $+\gamma_{2}$ and $-\gamma_{2}$ directions) to compute the signal and response. We use the GalSim package [^12] https://github.com/GalSim-developers/GalSim (Rowe et al., 2015) and the high-level routines from the ngmix package [^13] https://github.com/esheldon/ngmix for these image operation steps.
In previous DES shear analyses, the technique above was applied at the locations of objects detected in the main survey Gold catalogue. However, since object detection is shear-dependent especially in the presence of blending, Metacalibration causes a selection/detection bias which amounts to percent-level multiplicative bias in the shape measurement (see S20 for more detail). Instead, for Metadetection, we apply the Metacalibration artificial image shearing procedure to an entire cell-based coadd. This step generates five 200x200 cell coadds, one for each artificial shear, as described above. We then apply detection using the sep package (Barbary, 2016) to each cell-based coadd. We use the same settings as the main survey [^14] https://github.com/esheldon/sxdes. Finally, we compute the shear response by averaging over the catalogues of detections, after the same selections have been applied to all the catalogues, like so
@@
The difference between this procedure and Metacalibration is important. Metadetection cannot measure an individual object’s response. Further, because object detection depends on shear, one cannot match the five different sets of detections against one another [^15] The number of objects in each catalogue may differ.. The response above includes both the “shear” and “selection” responses used with the DES Y3 shear catalogues (Gatti, Sheldon et al., 2021). Finally, we note that separating sources into tomographic bins by photometric redshift is a selection cut that is done on the source magnitudes. Thus tomographic selection must be included in the selection cuts used to compute the response in eqn. 6, producing a different response estimate for each tomographic bin.
@@ 4.2.1 Pixel Masks
where $\mathrm{G}$ is the magnitude in the Gaia band. We further place a lower bound on the radius of 5 arcseconds. The image was set to zero inside the circular masked area, with the edges “apodized” to transition smoothly between unity on the outside to zero on the inside of the mask. This smooth transition was parameterized using the cumulative integral of a triweight kernel, which is a function of two parameters, $m$ and $h$, and is defined for a point $x$ with quantity $y=(x-m)/h$ as
@@ 4.2.2 Object Level Masks
    Metadetection footprint: An overall footprint for the initial Metadetection catalogue was produced by keeping areas corresponding to all cell-based coadds that had data in all four $griz$ bands and had no failures in running Metadetection. In this step, we retain the union of only the unique cell-based coadd and unique coadd tile regions as described in Sec. 3.
@@
    DES stars: We augmented the Gaia star catalogue with high-confidence DES stars, defined as MASH==0==0= = 0 and $r<21$ (Bechtol et al., 2025). These objects were masked with a fixed circular radius of 5 arcseconds. Many of these stars are also in Gaia  and thus may already have a larger mask applied.
@@
    HyperLEDA mask: After application of the foreground mask, we visually found additional unmasked regions near bright and extended objects at low redshift that caused spurious detections and contaminated the light profile of neighboring objects. We found that masking objects from the HyperLEDA (Paturel et al. 2003) significantly reduced spurious detections and poor measurements. The HyperLEDA catalogue contains galaxies brighter than $B$-mag $\sim 18$ with coordinates, diameter, and axis ratio. Each object was masked with a circular, $B$-magnitude dependent circular mask with a radius in degrees $0.147-0.00824*B$. Finally, for a handful of extremely large, low redshift objects, we identified spurious detections due to resolved structure within the object. Thus, we produced by hand special elliptical masks around those objects (NGC0055, NGC0253, NGC0300, IC1613) utilizing the diameter and axis ratio information of those objects in the HyperLEDA catalogue with some manual tuning.
@@ 4.3 Object Identification & Measurement
We detect and measure objects in the five different cell-based coadds Metadetection produces as follows. For detection, we use the sep package (Barbary, 2016), a Python-based fork of SExtractor (Bertin & Arnouts, 1996; Bertin, 2011). We use the same configuration setting as in DES collaboration et al. 2021 as contained in the sxdes [^17] https://github.com/esheldon/sxdes package. Given the coadd images for each band, we form the multi-band $riz$ coadd image through a straight addition of the coadds without any PSF matching, and run the detection on an inverse-variance weighted coadd image [^18] Early tests found that including the $g$-band in Metadetection reduced the precision of the shear estimates..
For each detection in each of the five images, we make three main measurements. All measurements were done with the ngmix software package. [^19] https://github.com/esheldon/ngmix First, we measure the average masked fraction per object. To make this measurement, we compute a single masked fraction image for the cell-based coadd across all of the bands used for lensing. In this work, we use an inverse-variance weighted average image of the cell-based coadd masked fraction images from $riz$ bands. Then for each detection, we compute the average value of this single masked fraction image with a two-arcsecond FWHM Gaussian weight function about the detection’s center. As for each object property measurement, we performed a Gaussian forward model fit across the $riz$ band images individually. Table 1 shows the list of parameters to fit each object and their priors.
@@
Total flux Two-side Erf min = $-1.0\times 10^{4}$, $\sigma_{\rm min}$ = 1.0,
@@
This fit follows a similar procedure as the DES Y3 Metacalibration shear catalogue (Gatti, Sheldon et al., 2021), including the use of uberseg (Jarvis et al., 2016) to remove some of the effects of light from neighboring objects. We refer to this measurement as the gauss quantities, and the shape from this measurement is used as the catalogue shape measurement with Metadetection. We additionally use a specialized Fourier-space method, detailed in Appendix B, to measure a two-arcsecond FWHM Gaussian weighted flux for each detection on a PSF-deconvolved image. This second set of measurements does not use uberseg. This second set of measurements is referred to as the pgauss quantities. These quantities include fluxes in each of the $griz$ bands along with an inverse variance-weighted size measure. Both gauss and pgauss measure pre-PSF properties of the object so that, e.g., point-like objects would have zero size in the absence of other effects.
While we do not detect with the $g$-band cell-based coadd, we use the detected object position to measure a $g$-band flux in addition to $riz$ fluxes. Measurements of the reconvolved PSF model of each band’s cell-based coadd are made with the pgauss method as well, but without PSF deconvolution. The default parameters and priors for all measurements are detailed in the metadetect [^20] https://github.com/esheldon/metadetect and pizza-cutter-metadetect [^21] https://github.com/beckermr/pizza-cutter-metadetect packages. The versions of key software packages used for each step of the pipeline are listed in Appendix A.
@@ 4.4.3 Flux/color & heavily interpolated objects
[…]
  • In order to reject objects that have odd colors that appear to be the result of flux measurement failures, we make cuts on measured galaxy color. $|g-r|<5$, $|r-i|<5$, $|i-z|<5$.
@@ 4.4.4 Junk detections
Figure 4: Log density of objects excluded by the star-galaxy separation criteria, showing object S/N vs size ratio ($T/T_{\rm PSF}$; left) space and color space ($g-r$ vs $r-i$; right). All other selection criteria have been applied. Objects are chosen from 100 randomly selected patches out of 200 over the footprint. Well-measured stars populate the plot near $T/T_{\rm PSF}=0$, but not exactly on $T/T_{\rm PSF}=0$, due to how the shape and size priors are set for the Gaussian fit.
Table 2: The fraction of objects removed with each criterion, computed as those in noshear Metadetection catalogue. We also show the fraction of stars within matched samples that are identified in Gold sample (MASH==0==0= = 0 OR $1$) within 0.263 arcseconds. Before “mask” selection, “basic” selections (flags==0==0= = 0 and $S/N^{\rm gauss}>5$) have been applied in order to save disk space.
@@ 4.5 Shear Catalogue Statistical Weights
The distributions of number counts, shape variance averaged in two components, shear response, and shear weight are shown in Fig. 5. We note here that the grid of shear response is smoothed with a Gaussian kernel of $\sigma=2.0$ to reduce the impact of shot noise of the response itself. We noticed objects with moderate S/N and large size having larger shape noise compared to Y3. We viewed those objects and found they are largely in a blended system. Nonetheless, the distribution of the shear weight is similar to that of the Metacalibration shape catalogue in Y3 (Gatti, Sheldon et al., 2021).
Figure 6: Left: weighted galaxy number density in units of gal/arcmin2 using Eq. 14. The number density is computed in a HEALPix grid (Górski et al., 2005; Zonca et al., 2019) of resolution 10 (nside$\,=2^{10}=1024$). Right: object shape noise computed with Eq. 13 in a HEALPix grid of resolution 10 (nside$\,=2^{7}=1024$).

### v2

@@ 2.3 Catalogue Blinding
[…]
@@ 3.1 Coadd Image Geometry
Figure 1: The cell-based coadd and DES coadd tile geometry for a corner of the DES coadd tile. The black dashed line shows the outer boundary of the $\times10k$ DES coadd tile. The inner black solid line shows the unique region of the coadd tile, which is defined on the sphere in right ascension and declination. Adjacent coadd tiles overlap the same area, but share adjacent unique region boundaries. The red solid lines show the unique regions for the cell-based coadds. In the interior of the coadd tile, these regions are $\times100$ pixels. The red dashed line (shown only for two of the cell-based coadds along the diagonal) shows the outer boundary of each cell-based coadd. Adjacent cell-based coadds with their outer boundaries overlap by 50 pixels. At the edge of the coadd tile, the unique region for the cell-based coadd is extended all the way to the outer boundary of the coadd tile.
@@ 3.2 Cell-based Coadd Construction and Data Products
Figure 2 shows example color images of nine adjacent cells from the DES Y6 analysis on the left, the mask in the center, and the masked fraction images on the right. We applied the cell-based coadding algorithm to the grizgriz bands and have produced a grigri false-color image for this figure. In this figure, we have cut the 50-pixel edge buffers and stitched the inner cell images together. We can see that there are subtle shifts in the background noise as the input SE images change from cell to cell. We also see some residual striping from the interpolation applied to the input SE images. This effect is accounted for using a cut in the masked fraction around each detection when running Metadetection, as described below. The masked fraction image shows that in regions where the interpolant is applied to nearly every input SE image, and thus the output cell-based coadd image is not usable, the masked fraction approaches unity. Any source detected in or near these highly masked regions will have a large masked fraction and is excluded from use in the catalogue. The bottom left cell in Figure 2 shows the effects of the 90-degree rotation applied to the pixel masks. This rotation has not been applied to all cells in the bottom row of the images because the cells on the right are marked as being near a bright star, and we do not rotate pixel masks near bright stars since that data will be excluded anyway in our final survey mask as discussed below.
Figure 2: Left: Example false-color image of nine coadd cells (with borders indicated as white dotted lines) made from grigri bands. The inner 100 $\times$ 100 pixels are cut out from the whole coadd cell of 200 $\times$ 200 pixels by removing the 50 pixels buffer region around the cell. Middle: The mask image over the same region of the sky, where the grey region is excluded from the detection and further analysis. Right: The mfrac image over the same region of the sky, averaged over rizriz-bands. For each detection object, the average masked fraction is computed from the mfrac coadd image, and any objects with mfrac > 0.1 are excluded from our analysis. The interpolation and masked fraction differs for the bottom left cell relative to the two other cells on the bottom row due to the fact that 90-degree rotations are not applied to the input pixel masks near bright stars. See Sec. 3 for more details.
@@ 4.1 Shear Calibration Formalism
Metacalibration (Huff and Mandelbaum 2017; Sheldon and Huff 2017) was the first algorithm to implement this concept into practice to self-calibrate “raw” galaxy shapes. Metacalibration operates on existing object detections and includes corrections for non-detection dependent selection effects implicitly in the ensemble average in Eqn. 3. This technique deconvolves the PSF from the image, applies a small artificial shear, and then reconvolves the image with a slightly larger PSF. To this artificially sheared image, we add a noise image put through the same process but with an opposite shear applied. [^11] In detail, we rotate the noise image by 90 degrees, run an identical shearing process as the original image, and then rotate the noise image back by 90 degrees.. This correction accounts for the effect of sheared and correlated background noise on the shape measurement (Sheldon and Huff, 2017). Without it, the technique will produce catastrophic biases at the order of $\sim 10\%$. Finally, from the measured shapes on artificially sheared images, we can use a two-sided finite difference formula to compute the shear response
@@
where the subscript $i$ denotes one of the two shear components and the superscript +/- denotes the sign of the applied shear. In total, we create five images (unsheared, sheared in $\gamma_{1},-\gamma_{1}$, $\gamma_{2}$ and $\gamma_{2}$ directions with the amplitude of 0.01) to compute the signal and response. We use the GalSim package [^12] https://github.com/GalSim-developers/GalSim (Rowe et al., 2015) and the high-level routines from the ngmix package [^13] https://github.com/esheldon/ngmix for these image operation steps.
In previous DES shear analyses, the technique above was applied at the locations of objects detected in the main survey Gold catalogue. However, since object detection is shear-dependent especially in the presence of blending, Metacalibration causes a selection/detection bias which amounts to percent-level multiplicative bias in the shape measurement (see S20 for more detail). Instead, for Metadetection, we apply the Metacalibration artificial image shearing procedure to an entire cell-based coadd. This step generates five 200 $\times$ 200 cell coadds, one for each artificial shear, as described above. We then apply detection using the sep package (Barbary, 2016) to each cell-based coadd. We use the same settings as the main survey [^14] https://github.com/esheldon/sxdes. Finally, we compute the shear response by averaging over the catalogues of detections, after the same selections have been applied to all the catalogues, like so
@@
The difference between this procedure and Metacalibration is important. Metadetection cannot measure an individual object’s response. Further, because object detection depends on shear, one cannot match the five different sets of detections against one another [^15] The number of objects in each catalogue may differ.. The response above includes both the “shear” and “selection” responses used with the DES Y3 shear catalogues (Gatti, Sheldon, and others, 2021). Finally, we note that separating sources into tomographic bins by photometric redshift is a selection cut that is done on the source magnitudes. Thus tomographic selection must be included in the selection cuts used to compute the response in eqn. 6, producing a different response estimate for each tomographic bin.
@@ 4.2.1 Pixel Masks
where $\mathrm{G}$ is the magnitude in the Gaia band. We further place a lower bound on the radius of 5 arcseconds. The image was set to zero inside the circular masked area, with the edges “apodized” to transition smoothly between unity on the outside to zero on the inside of the mask. This smooth transition was parameterized using the cumulative integral of a triweight kernel, which is a function of two parameters, $m$ and $h$, and is defined for a point $x$ with quantity y=(xm)/hy=(x-m)/h as
@@ 4.2.2 Object Level Masks
    Metadetection footprint: An overall footprint for the initial Metadetection catalogue was produced by keeping areas corresponding to all cell-based coadds that had data in all four grizgriz bands and had no failures in running Metadetection. In this step, we retain the union of only the unique cell-based coadd and unique coadd tile regions as described in Sec. 3.
@@
    DES stars: We augmented the Gaia star catalogue with high-confidence DES stars, defined as MASH==0==0 and $r<21$ (Bechtol et al., 2025). These objects were masked with a fixed circular radius of 5 arcseconds. Many of these stars are also in Gaia  and thus may already have a larger mask applied.
@@
    HyperLEDA mask: After application of the foreground mask, we visually found additional unmasked regions near bright and extended objects at low redshift that caused spurious detections and contaminated the light profile of neighboring objects. We found that masking objects from the HyperLEDA (Paturel et al. 2003) significantly reduced spurious detections and poor measurements. The HyperLEDA catalogue contains galaxies brighter than $B$-mag $\sim 18$ with coordinates, diameter, and axis ratio. Each object was masked with a circular, $B$-magnitude dependent circular mask with a radius in degrees 0.1470.00824B0.147-0.00824*B. Finally, for a handful of extremely large, low redshift objects, we identified spurious detections due to resolved structure within the object. Thus, we produced by hand special elliptical masks around those objects (NGC0055, NGC0253, NGC0300, IC1613) utilizing the diameter and axis ratio information of those objects in the HyperLEDA catalogue with some manual tuning.
@@ 4.3 Object Identification & Measurement
We detect and measure objects in the five different cell-based coadds Metadetection produces as follows. For detection, we use the sep package (Barbary, 2016), a Python-based fork of SExtractor (Bertin and Arnouts, 1996; Bertin, 2011). We use the same configuration setting as in DES collaboration et al. 2021 as contained in the sxdes [^17] https://github.com/esheldon/sxdes package. Given the coadd images for each band, we form the multi-band rizriz coadd image through a straight addition of the coadds without any PSF matching, and run the detection on an inverse-variance weighted coadd image [^18] Early tests found that including the $g$-band in Metadetection reduced the precision of the shear estimates..
For each detection in each of the five images, we make three main measurements. All measurements were done with the ngmix software package. [^19] https://github.com/esheldon/ngmix First, we measure the average masked fraction per object. To make this measurement, we compute a single masked fraction image for the cell-based coadd across all of the bands used for lensing. In this work, we use an inverse-variance weighted average image of the cell-based coadd masked fraction images from rizriz bands. Then for each detection, we compute the average value of this single masked fraction image with a two-arcsecond FWHM Gaussian weight function about the detection’s center. As for each object property measurement, we performed a Gaussian forward model fit across the rizriz band images individually. Table 1 shows the list of parameters to fit each object and their priors.
@@
Total flux Two-side Erf min = $1.0\times 10^{4}$, $\sigma_{\rm min}$ = 1.0,
@@
This fit follows a similar procedure as the DES Y3 Metacalibration shear catalogue (Gatti, Sheldon, and others, 2021), including the use of uberseg (Jarvis et al., 2016) to remove some of the effects of light from neighboring objects. We refer to this measurement as the gauss quantities, and the shape from this measurement is used as the catalogue shape measurement with Metadetection. We additionally use a specialized Fourier-space method, detailed in Appendix B, to measure a two-arcsecond FWHM Gaussian weighted flux for each detection on a PSF-deconvolved image. This second set of measurements does not use uberseg. This second set of measurements is referred to as the pgauss quantities. These quantities include fluxes in each of the grizgriz bands along with an inverse variance-weighted size measure. Both gauss and pgauss measure pre-PSF properties of the object so that, e.g., point-like objects would have zero size in the absence of other effects.
While we do not detect with the $g$-band cell-based coadd, we use the detected object position to measure a $g$-band flux in addition to rizriz fluxes. Measurements of the reconvolved PSF model of each band’s cell-based coadd are made with the pgauss method as well, but without PSF deconvolution. The default parameters and priors for all measurements are detailed in the metadetect [^20] https://github.com/esheldon/metadetect and pizza-cutter-metadetect [^21] https://github.com/beckermr/pizza-cutter-metadetect packages. The versions of key software packages used for each step of the pipeline are listed in Appendix A.
@@ 4.4.3 Flux/color & heavily interpolated objects
  • In order to obtain reliable photometric redshifts of the sources, we make magnitude [^22] Magnitudes were estimated from the pgauss fluxes with an inverse hyperbolic sine function (Lupton et al., 1999). The magnitudes were further corrected for the extinction using E(BV)E(B-V) values from the reddening map of Schlegel et al. (1998), which is at the HEALPix resolution of nside=4096. cuts on each bandpass to exclude excessively faint sources following Myles, Alarcon, and others (2021). These cuts are $g<26.5$, $r<26.5$, i<24.7i<24.7232323This selection is to reduce COSMOS20 photometric redshift outliers (Weaver et al., 2022) with which our redshift calibration is anchored (Yin and others, 2025). , and $z<25.6$.
  • In order to reject objects that have odd colors that appear to be the result of flux measurement failures, we make cuts on measured galaxy color. |gr|<5|g-r|<5, |ri|<5|r-i|<5, |iz|<5|i-z|<5.
@@ 4.4.4 Junk detections
Figure 4: Log density of objects excluded by the star-galaxy separation criteria, showing object S/N vs size ratio ($T/T_{\rm PSF}$; left) space and color space (grg-r vs rir-i; right). All other selection criteria have been applied. Objects are chosen from 100 randomly selected patches out of 200 over the footprint. Well-measured stars populate the plot near $T/T_{\rm PSF}=0$, but not exactly on $T/T_{\rm PSF}=0$, due to how the shape and size priors are set for the Gaussian fit.
Table 2: The fraction of objects removed with each criterion, computed as those in noshear Metadetection catalogue. We also show the fraction of stars within matched samples that are identified in Gold sample (MASH==0==0 OR 11) within 0.263 arcseconds. Before “mask” selection, “basic” selections (flags==0==0 and $S/N^{\rm gauss}>5$) have been applied in order to save disk space.
@@ 4.5 Shear Catalogue Statistical Weights
The distributions of number counts, shape variance averaged in two components, shear response, and shear weight are shown in Fig. 5. We note here that the grid of shear response is smoothed with a Gaussian kernel of $\sigma=2.0$ to reduce the impact of shot noise of the response itself. We noticed objects with moderate S/N and large size having larger shape noise compared to Y3. We viewed those objects and found they are largely in a blended system. Nonetheless, the distribution of the shear weight is similar to that of the Metacalibration shape catalogue in Y3 (Gatti, Sheldon, and others, 2021).
[…]